It has been reported that 70% of university students do volunteer work during their summer vacation. Four students are randomly selected to do volunteer work.


a. The probability that at least 1 student will do volunteer work this summer (correct to 3 decimal places) is 


b. The probability that exactly 3 graduates will not do any volunteer work this summer (correct to 4 decimal places) is 


c. The expected number of students (correct to 1 decimal place) who will not do volunteer work this summer is


1
Expert's answer
2022-05-11T12:46:06-0400

Let "X=" the number of students who do volunteer work during their summer vacation:"X\sim Bin(n, p)."

Given "n=4, p=0.7, q=1-p=0.3."

a.


"P(X\ge1)=1-P(X=0)"

"=1-\dbinom{4}{0}(0.7)^0(0.3)^{4-0}=0.992"

b.


"P(X^C=3)=P(X=1)=\dbinom{4}{1}(0.7)^1(0.3)^{4-1}"

"=0.0756"

c.

"E(X^C)=nq=4(0.3)=1.2"